A 2700 ohm resistor and a 1.1uF capacitor are connected in series across a generator (60 Hz, 120 V). Determine the average power consumed in the circuit. a.) 3W b.) 9W c.) 10W d.) 7W

Grade:

1 Answers

Arun
5 years ago
Dear Sarthak
 
X_C = 1/(2π f C) = 1/(2π 60Hz * 1.1 * 10^[-6]) = 2411Ω 

Z = √ (R^2 + X_C^2) = √ ([2700Ω]^2 + [2411Ω]^2) = 3620Ω 

I = V_S/Z = 120V/3620Ω = 33.1mA 

Now there are three powers in this circuit. Real power across resistor, reactive power across capacitor and apparent power source delivers. 

Real P = I^2 R = (33.1mA)^2 * 2700Ω = 2.97W --- this is probably your answer 
Reactive Q_C = I^2 X_C = (33.1mA)^2 * 2411Ω = 2.65VAR 
Apparent S = I^2 Z = (33.1mA)^2 * 3620Ω = 3.98V•A
 
Regards
Arun (askIITians forum expert)